OpenStudy (anonymous):

Reviewing my algebra please check. What are the zeros of \(y = x^2 - 2x -15\)? (-3, 5) is what I got.

OpenStudy (bibby):

0=(x-5)(x+3) x=5,3

OpenStudy (anonymous):

Set the equation to equal to zero: $$x^2 - 2x -15 = 0$$ $$(x - 5)(x + 3)$$ Use FOIL: $$x^2 + 3x - 5x - 15$$ $$x^2 - 2x - 15$$ $$(5 - 5) = 0 = (-3 + 3)$$ \(\textbf{Hence the zeros are:}\) (-3, 5) So you are correct.

OpenStudy (anonymous):

Well, hold on guys. Don't out them in (-3, 5) format

OpenStudy (anonymous):

Are there any other methods I could use other than this?

OpenStudy (anonymous):

That's a coordinate, tecchnically, and would be marked wrong by a teacherif placed as the answer

OpenStudy (bibby):

that isn't meant to be a non-inclusive range but rather the 2 discrete values

OpenStudy (anonymous):

I believe graphing is another method. Not 100% sure.

OpenStudy (anonymous):

You could do that.

OpenStudy (anonymous):

@LanguageEnthusiast want me to show you?

OpenStudy (anonymous):

@Algorithmic sure if you don't mind :3 I would hate to be a bother.

OpenStudy (anonymous):

Imagine a graph as such:  See where I circled the graph, the place where the parabola intersects the x-axis are where the zeros would be.

OpenStudy (anonymous):

Oh I see! I remember this actually thank you so much! :-)

OpenStudy (anonymous):

You are welcome, also it is never a bother in my opinion to ask for knowledge. So ask away. :-)